OpenStudy (anonymous):

How to solve integral f(u) = e^(u/u-1)*[1/(1-u)^2] ? I will retype this clear in my reply. Thanks

OpenStudy (anonymous):

\[\int\limits_{-∞}^{+∞} e^{\frac{ u }{ u-1 }} * \frac{ 1 }{ (1-u)^2 }\]

OpenStudy (dumbcow):

hmm lets try a substitution... \[w = \frac{u}{u-1}\] \[dw = \frac{-1}{(u-1)^{2}} du\] note: \[(u-1)^{2} = (1-u)^{2}\] \[\rightarrow - \int\limits e^{w} dw\]

OpenStudy (anonymous):

so, which means it equal to \[-e^{w} + C\] ?

OpenStudy (dumbcow):

yes...sub back in to get in terms of "u"

OpenStudy (anonymous):

Thanks @dumbcow may I ask in further if 0<u<1, means w also 0<w<1 so is that means \[-\int\limits_{0}^{1} e^{w} = 1?\]

OpenStudy (dumbcow):

no the limits change whenever you make a substitution....thats why after integrating you sub back in the original variable you can also change the limits but i prefer not to

OpenStudy (zzr0ck3r):

same here

OpenStudy (dumbcow):

\[-\int\limits e^{w} = -e^{w} = -e^{u/(u-1)}|_0^1\]

OpenStudy (anonymous):

so \[-e^{1/0} + e^{0} = \infty + 1 ?\]

OpenStudy (dumbcow):

\[-\infty +1\] :)

OpenStudy (anonymous):

Thanks :)